# Thermally driven two-sphere microswimmer with internal feedback control

### v1

@@ Abstract
We discuss the locomotion of a thermally driven elastic two-sphere microswimmer with internal feedback control that is realized by the position-dependent friction coefficients. In our model, the two spheres are in equilibrium with independent heat baths having different temperatures, causing a heat flow between the two spheres. We generally show that the average velocity of the microswimmer is nonzero when the friction coefficients are position-dependent. Using the method of stochastic thermodynamics, we obtain the entropy production rate and discuss the efficiency of the two-sphere microswimmer. The proposed self-propulsion mechanism highlights the importance of information in active matter and can be a fundamental process in various biological systems.
@@ I Introduction
Among various generalizations of the three-sphere microswimmer model Yasuda23 , Hosaka et al. considered an elastic three-sphere microswimmer in which the three spheres are in equilibrium with independent heat baths having different temperatures Hosaka17 . It was shown that such a stochastic microswimmer (without any prescribed deformation) can also acquire net locomotion due to thermal fluctuations, and its average velocity is proportional to the heat flow among the spheres Hosaka17 . Later, the average entropy production rate Sou19 ; Sou21 and the time-correlation functions Li24 of the same model were calculated by some of the current authors. In particular, the existence of an antisymmetric part of the cross-correlation function reflects the broken time-reversal symmetry of this microswimmer Li24 .
@@
Prior to the above thermal three-sphere microswimmer model, Kumar et al. proposed a model of an active elastic dimer (AED) in which the friction coefficients of the two spheres depend on their relative coordinate Kumar08 ; Baule08 . Their model is different from the traditional Brownian ratchet models Julicher97 ; Reimann because the motion asymmetry is created internally and is not induced by an external periodic potential. They found that the average velocity of an AED is proportional to the difference between the non-equilibrium noise strengths acting on the two particles Kumar08 ; Baule08 . Such a self-propulsion mechanism can explain, for example, the movement of helicases on DNA Yu06 , the walking of Myosin VI on actin filaments Altman04 , and the collective migration of cell clusters Pages22 .
@@
Given the modern framework of stochastic thermodynamics SekimotoBook ; Jarzynski11 ; Seifert12 and the accumulated knowledge of the thermal three-sphere microswimmer model Hosaka17 ; Sou19 ; Sou21 , the AED model offers a renewed interest, especially when the two spheres have different temperatures without any non-equilibrium noises. Moreover, the internal feedback control considered in the AED model is a crucial mechanism for “informational active matter” that utilizes information (measurement and feedback) instead of energy for various non-equilibrium processes VanSaders24 ; Huang20 . Recently, we have proposed models of Ornstein-Uhlenbeck information swimmers with external and internal feedback controls Hou24 .
@@
In this work, we simplify the original AED model to discuss a thermally driven two-sphere microswimmer with internal feedback control that is realized by the position-dependent friction coefficients. In the current model, the two spheres are in equilibrium with independent heat baths having different temperatures. We show that a combination of heat transfer between the spheres and internal feedback control leads to directional locomotion in the steady state under a noisy condition. We analytically obtain the average velocity and the entropy production rate of this stochastic microswimmer by using stochastic energetics SekimotoBook . With these results, we further estimate the efficiency of the thermal two-sphere microswimmer. We consider that our model explains the fundamental mechanism for transport and locomotion in various active matter and biological systems.
@@
In Sec. II, we explain the model of a thermally driven elastic two-sphere microswimmer with internal feedback control. In Sec. III, we obtain the steady-state probability distribution that is further used to calculate the average velocity and the entropy production rate in Secs. IV and V, respectively. In Sec. VI, we calculate the efficiency of the two-sphere microswimmer by using specific functional forms for the spring potential and the friction coefficients. A brief discussion and summary is given in Sec. VII.
@@
Figure 1: Thermally driven elastic microswimmer consisting of two hard spheres connected by a spring. The positions of the two spheres in a one-dimensional coordinate are denoted by $x_{i}(t)$ ($i=1,2$), and the spring extension is defined by $x=x_{2}-x_{1}$. The spring potential energy is given by $U(x)$, and the two spheres are in equilibrium with independent heat baths having different temperatures $T_{i}$. The friction coefficients $\gamma_{i}(x)$, acting separately on the two spheres, are dependent on the spring extension $x$.
@@ II Model
As schematically shown in Fig. 1, we consider a microswimmer consisting of two hard spheres connected by a spring. The positions of the two spheres in a one-dimensional coordinate are denoted by $x_{i}(t)$ ($i=1,2$), and we further define the spring extension by $x=x_{2}-x_{1}$ that can take both positive and negative values. The spring potential energy is generally given by $U(x)$, which can be a harmonic potential (as we will discuss later) or that of a finitely extensible nonlinear elastic (FENE) spring Baule08 . There are two important assumptions of the model: (i) the two spheres are in equilibrium with independent heat baths having different temperatures $T_{i}$ ($T_{1}\neq T_{2}$) and (ii) the friction coefficients $\gamma_{i}(x)\geq 0$, acting separately on the two spheres, are dependent on the spring extension $x$. To keep the generality of the model as much as possible, we first do not specify the functional forms of $U(x)$ and $\gamma_{i}(x)$.
@@
where $\dot{x}_{i}=dx_{i}/dt$, $\partial_{i}=\partial/\partial x_{i}$ (the summation over $i$ is not taken). The thermal noise $\xi_{i}(t)$ has Gaussian statistics with zero mean and unit variance:
@@
According to the fluctuation-dissipation relation, the prefactor of thermal noise is given by $g_{i}(x)=\sqrt{2\gamma_{i}(x)k_{\rm B}T_{i}}$, where $k_{\rm B}$ is the Boltzmann constant DoiBook . The above Langevin equations are multiplicative because the noise is coupled nonlinearly with the stochastic variables Kampenbook . Then, it is necessary to specify the interpretation of the multiplicative noise; otherwise Eq. (1) is meaningless.
Here, we employ the Itô interpretation Gardinerbook , and the Langevin equations in Eq. (1) should be rewritten with the Itô convention $x^{\ast}=x(t)$ as
@@
As pointed out by Lau and Lubensky Lau07 , the second term on the right-hand side is proportional to the temperature, and it is necessary to ensure the proper thermal equilibrium. Such a term reflects thermal fluctuation and generally exists for other interpretations except for the anti-Itô convention Kuroiwa14 . Hereafter, the prime denotes a derivative with respect to the relative coordinate $x=x_{2}-x_{1}$, such as $U^{\prime}(x)=-\partial_{1}U(x)=\partial_{2}U(x)$. Then, the coupled Langevin equations can be explicitly written as
@@ III Steady-state probability distribution
From Eqs. (4) and (5), the equations of motion for both the center-of-mass coordinate $X=(x_{1}+x_{2})/2$ and the relative coordinate $x=x_{2}-x_{1}$ can be obtained. Since the former equation is a function of the relative coordinate $x$ and the noise only, we first discuss the equation of motion for $x$ in the Itô interpretation given by Baule08 .
@@
where we have introduced the temperature ratio $\theta=T_{2}/T_{1}$, i.e., $\theta=1$ corresponds to the thermal equilibrium.
The Fokker-Planck equation for the probability distribution $P(x,t)$ associated with the Langevin equation in Eq. (6) can be written in the Itô interpretation as Riskenbook
@@
Since we are interested in the steady-state properties of the microswimmer, we obtain the steady-state probability distribution by imposing $\partial_{t}P(x,t)=0$. Moreover, the probability flux can be set to zero since the relative coordinate is bounded by the spring potential $U(x)$. Then, the steady-state probability distribution $P_{\rm s}(x)$ satisfies Baule08
@@
which is essentially the Boltzmann distribution modified by the two different temperatures.
@@ IV Average velocity
Having obtained the steady-state distribution function $P_{\mathrm{s}}(x)$ for the relative coordinate $x$, we can calculate the average velocity of the elastic two-sphere microswimmer. Since the center of mass velocity is simply given by $V=\dot{X}=(\dot{x}_{1}+\dot{x}_{2})/2$, we use Eqs. (4) and (5) to obtain the statistical average of $V$ Kumar08 ; Baule08
@@
This result clearly demonstrates that the microswimmer acquires a finite average velocity when $\theta\neq 1$ ($T_{1}\neq T_{2}$) and $\gamma_{1}$ is not constant. In other words, the current microswimmer is driven by thermal energy, and the internal feedback control through $\gamma_{1}(x)$ plays an essential role in the locomotion. However, when $(1/\gamma_{1}(x))^{\prime}$ and $U(x)$ are respectively odd and even functions of $x$, $\langle V\rangle$ vanishes even if the friction coefficient is position-dependent. A similar result to Eqs. (14) and (15) was reported for an AED Kumar08 ; Baule08 , for which it was shown that an asymmetry in the non-equilibrium noise strengths leads to locomotion. In our study, we do not introduce any active fluctuations, which allows us to use the standard framework of stochastic thermodynamics SekimotoBook ; Jarzynski11 ; Seifert12 , as we discuss below.
Figure 2: The average velocity [see Eq. (22)] and the average entropy production rate [see Eq. (23)] of the two-sphere microswimmer when the potential is $U(x)=Kx^{2}/2$ and the friction coefficients are $\gamma_{1}(x)=\gamma_{2}(x)=\gamma[1+\delta\tanh(x/w)]$. Here, we choose $\varepsilon=Kw^{2}/(k_{\mathrm{B}}T_{1})=1$. The filled circles correspond to the result of the numerical simulation. (a) The plot of the dimensionless average velocity $\langle\hat{V}\rangle=\langle V\rangle\gamma/\sqrt{Kk_{\mathrm{B}}T_{1}}$ as a function of $\delta$ for different values of $\theta=T_{2}/T_{1}$. (b) The plot of the dimensionless average entropy production rate $\langle\hat{\dot{\sigma}}\rangle=\langle\dot{\sigma}\rangle\gamma T_{1}/(Kw)^{2}$ as a function of $\delta$ for different values of $\theta$.
Figure 3: The color plot of the efficiency $\eta$ [see Eq. (24)] of the two-sphere microswimmer as a function of $\delta$ and $\theta$. The efficiency $\eta$ vanishes when $\delta=0$.
@@ V Entropy production rate
Next, we calculate the average entropy production rate of the elastic microswimmer in the steady state, which is given by
@@
where the stochastic heat flow $\dot{Q}_{i}$ is the heat gained by the $i$-th sphere per unit time. According to the stochastic energetics SekimotoBook , it is given by
@@
where $\circ$ denotes the Stratonovich product, and the small increment $dx_{i}$ can be obtained from the Langevin equations in Eq. (3). Using the Wong-Zakai theorem to deal with the Stratonovich product and taking the statistical average SekimotoBook , we obtain after some calculation
@@
One can easily check that $\langle\dot{Q}_{i}\rangle=0$ when $\theta=1$ by using the equilibrium distribution function in Eq. (12). When $\theta\neq 1$, we consider the total heat flow defined by $\langle\dot{Q}\rangle=\langle\dot{Q}_{1}\rangle+\langle\dot{Q}_{2}\rangle$. By using Eqs. (7) and (8), one can generally show that $\langle\dot{Q}\rangle$ vanishes in the steady state because
@@
In the second equality, we have used Eq. (10), which comes from the steady-state condition. Notice that the relation $\langle\dot{Q}\rangle=0$ generally holds even if the two friction coefficients $\gamma_{1}(x)$ and $\gamma_{2}(x)$ are different. Since we have $\langle\dot{Q}_{2}\rangle=-\langle\dot{Q}_{1}\rangle$ from Eq. (19), the average entropy production rate in Eq. (16) can be rewritten as

### v2

@@ Abstract
We investigate the locomotion of a thermally driven elastic two-sphere microswimmer with internal feedback control that is realized by the position-dependent friction coefficients. In our model, the two spheres are in equilibrium with independent heat baths characterized by different temperatures, causing a heat flow between the two spheres. We generally show that the average velocity of the microswimmer is non-zero when the friction coefficients are dependent on the spring extension. Using the method of stochastic thermodynamics, we obtain the entropy production rate and estimate the efficiency of the two-sphere microswimmer. The proposed self-propulsion model highlights the importance of information in active matter and offers a fundamental transport mechanism in various biological systems.
@@ I Introduction
Among various generalizations of the three-sphere microswimmer model Yasuda23 , Hosaka et al. proposed an elastic three-sphere microswimmer in which the three spheres are in equilibrium with independent heat baths having different temperatures Hosaka17 . It was shown that such a stochastic microswimmer (without any prescribed deformation) can also acquire net locomotion due to thermal fluctuations, and its average velocity is proportional to the heat flow between the spheres Hosaka17 . Later, the average entropy production rate Sou19 ; Sou21 and the time-correlation functions Li24 of the same model were calculated by some of the current authors. In particular, the existence of an antisymmetric part of the cross-correlation function reflects the broken time-reversal symmetry of this microswimmer Li24 .
@@
Prior to the above thermal three-sphere microswimmer model, Kumar et al. proposed a model of an active elastic dimer (AED) in which the friction coefficients of the two spheres depend on their relative coordinate Kumar08 ; Baule08 . Their model is different from the traditional Brownian ratchet models Julicher97 ; Reimann because the motion asymmetry is created internally and is not induced by an external periodic potential. They found that the average velocity of an AED is proportional to the difference between the non-equilibrium noise strengths acting on the two spheres Kumar08 ; Baule08 . Such a self-propulsion mechanism can explain, for example, the movement of helicases on DNA Yu06 , the walking of Myosin VI on actin filaments Altman04 , and the collective migration of cell clusters Pages22 .
@@
Given the modern framework of stochastic thermodynamics SekimotoBook ; Jarzynski11 ; Seifert12 and the accumulated knowledge of the thermal three-sphere microswimmer model Hosaka17 ; Sou19 ; Sou21 ; Li24 , the AED model offers a renewed interest, especially when the two spheres have different temperatures without any non-equilibrium noises. Moreover, the internal feedback control through the position-dependent friction coefficients in the AED model is a crucial mechanism for “informational active matter” that utilizes information (measurement and feedback) instead of energy for various non-equilibrium processes VanSaders24 ; Huang20 . Recently, we have proposed models of Ornstein-Uhlenbeck information swimmers with external and internal feedback controls Hou24 .
@@
In this work, we simplify the original AED model to discuss a thermally driven two-sphere microswimmer with internal feedback control that is realized by the position-dependent friction coefficients. In the current model, the two spheres are in thermal equilibrium with independent heat baths having different temperatures. We show that a combination of heat transfer between the spheres and internal feedback control leads to directional locomotion in a steady state under a noisy environment. We analytically obtain the average velocity and the entropy production rate of this stochastic micromachine by using stochastic energetics SekimotoBook . Using these results, we further estimate the efficiency of the proposed model. We consider that our model explains one of the fundamental mechanisms for transport and locomotion in various active matter and biological systems, such as the crawling motion of a cell Leoni17 ; Tarama18 .
@@
In Sec. II, we explain the model of a thermally driven elastic two-sphere microswimmer with internal feedback control. In Sec. III, we obtain the steady-state probability distribution that is further used to calculate the average velocity and the entropy production rate in Secs. IV and V, respectively. In Sec. VI, we estimate the efficiency of the two-sphere microswimmer by using specific functional forms for the spring potential and the friction coefficients. A brief summary and discussion are given in Sec. VII.
@@
Figure 1: (Color online) Thermally driven elastic microswimmer consisting of two hard spheres connected by a spring. The positions of the two spheres in a one-dimensional coordinate are denoted by $x_{i}(t)$ ($i=1,2$), and the spring extension is defined by $x=x_{2}-x_{1}$. The spring potential energy is given by $U(x)$, and the two spheres are in thermal equilibrium with independent heat baths having different temperatures $T_{i}$. The friction coefficients $\gamma_{i}(x)$, acting separately on the two spheres, are dependent on the spring extension $x$. The internal heat flow combined with the feedback control through the position-dependent friction coefficient leads to persistent locomotion of the microswimmer.
@@ II Model
As schematically shown in Fig. 1, we consider a microswimmer consisting of two hard spheres connected by a spring. The positions of the two spheres in a one-dimensional coordinate are denoted by $x_{i}(t)$ ($i=1,2$), and we further define the spring extension by $x=x_{2}-x_{1}$ that can take both positive and negative values. The spring potential energy is generally given by $U(x)$, which can be a harmonic potential (as we will discuss later) or that of a finitely extensible non-linear elastic (FENE) spring Baule08 . There are two important assumptions of the model: (i) the two spheres are in equilibrium with independent heat baths having different temperatures $T_{i}$ ($T_{1}\neq T_{2}$) and (ii) the friction coefficients $\gamma_{i}(x)\geq 0$, acting separately on the two spheres, are dependent on the spring extension $x$. The fact that the friction coefficients $\gamma_{i}(x)$ depend on the spring extension $x$ reflects an internal feedback mechanism: as the relative position of the spheres changes, the resistance experienced by each sphere varies accordingly. This coupling allows the system to modulate its dissipative response based on internal configuration and is essential for achieving directional locomotion in the absence of external forcing. To preserve the generality of the model, we first do not specify the functional forms of $U(x)$ and $\gamma_{i}(x)$.
@@
where $\dot{x}_{i}=dx_{i}/dt$, $\partial_{i}=\partial/\partial x_{i}$ (the summation over $i$ is not taken here). The thermal noise $\xi_{i}(t)$ has Gaussian statistics with zero mean and unit variance:
@@
According to the fluctuation-dissipation relation of the second kind, the prefactor of thermal noise is given by $g_{i}(x)=\sqrt{2\gamma_{i}(x)k_{\rm B}T_{i}}$, where $k_{\rm B}$ is the Boltzmann constant DoiBook . The above Langevin equations are multiplicative because the noise is coupled non-linearly with the stochastic variables Kampenbook . Then, it is necessary to specify the interpretation of the multiplicative noise (otherwise, Eq. (1) is meaningless).
Here, we employ the Itô interpretation Gardinerbook , and the Langevin equation in Eq. (1) should be rewritten with the Itô convention $x^{\ast}=x(t)$ as Baule08
@@
As pointed out by Lau and Lubensky Lau07 , the second term on the right-hand side is proportional to the temperature $T_{i}$, and it is necessary to ensure the proper thermal equilibrium. In principle, one can also choose other interpretations of the multiplicative noise, such as the Stratonovich interpretation Gardinerbook , as long as the same thermal equilibrium is ensured. Using, for example, the Stratonovich convention $x^{\ast\ast}=[x(t+\Delta t)+x(t)]/2$, where $\Delta t$ is the small increment of time, the corresponding Langevin equation should read (see the Appendix of Ref. Baule08 )
$\dot{x}_{i}(t)=-\dfrac{\partial_{i}U(x)}{\gamma_{i}(x)}-\dfrac{g_{i}(x)\partial_{i}g_{i}(x)}{2[\gamma_{i}(x)]^{2}}+\dfrac{g_{i}(x^{\ast\ast})}{\gamma_{i}(x^{\ast\ast})}\xi_{i}(t).$ (4)
Notice that Eqs. (3) and (4) statistically describe the same physical phenomena experiencing a multiplicative noise, and the second terms on the right-hand side of these equations vanish when the noise is additive, namely, when $\gamma_{i}(x)$ is constant. A more general argument, including the anti-Itô convention, can be seen in the literatures Lau07 ; Kuroiwa14 .
For the sake of mathematical convenience, we shall use the Itô interpretation and deal with the coupled Langevin equations in Eq. (3) that can be explicitly written as
@@
Hereafter, the prime denotes a derivative with respect to the relative coordinate $x=x_{2}-x_{1}$, such as $U^{\prime}(x)=-\partial_{1}U(x)=\partial_{2}U(x)$.
@@ III Steady-State Probability Distribution
From the Langevin equations in Eqs. (5) and (6), the equations of motion for both the center-of-mass coordinate $X=(x_{1}+x_{2})/2$ and the relative coordinate $x=x_{2}-x_{1}$ can be obtained. Since the former equation is a function of the relative coordinate $x$ and the noise only, we first discuss the equation of motion for $x$ within the Itô interpretation, which is given by Baule08
@@
where we have introduced the dimensionless temperature ratio $\theta=T_{2}/T_{1}$. Note that $\theta=1$ corresponds to the thermal equilibrium.
The Fokker-Planck equation for the probability distribution $P(x,t)$ corresponding to the Langevin equation in Eq. (7) can be written in the Itô interpretation as Riskenbook
@@
Since we are interested in the steady-state properties of the microswimmer, we discuss the steady-state probability distribution by imposing $\partial_{t}P(x,t)=0$. Moreover, the probability flux can be set to zero since the relative coordinate is bounded by the spring potential $U(x)$. Then, the steady-state probability distribution $P_{\rm s}(x)$ satisfies the relation Baule08
@@
which is essentially the Boltzmann distribution modified by the temperature ratio $\theta=T_{2}/T_{1}$.
@@ IV Average Velocity
Having obtained the steady-state distribution function $P_{\mathrm{s}}(x)$ for the relative coordinate $x$ in the previous section, we calculate the average velocity of the elastic two-sphere microswimmer. Since the center of mass velocity is simply given by $V=\dot{X}=(\dot{x}_{1}+\dot{x}_{2})/2$, we use Eqs. (5) and (6) to obtain the statistical average of $V$ Kumar08 ; Baule08
@@
This result clearly demonstrates that the microswimmer acquires a finite average velocity when $\theta\neq 1$ ($T_{1}\neq T_{2}$) and $\gamma_{1}$ is not constant. In other words, the current microswimmer is driven by thermal energy, and the internal feedback control through $\gamma_{1}(x)$ plays a crucial role for its locomotion. We remark, however, that when $(1/\gamma_{1}(x))^{\prime}$ and $U(x)$ are respectively odd and even functions of $x$, $\langle V\rangle$ vanishes even if the friction coefficient is position-dependent. A similar result to Eqs. (15) and (16) was reported for an AED whose asymmetry in the non-equilibrium noise strengths leads to locomotion Kumar08 ; Baule08 . In our study, we do not introduce any active fluctuations, which allows us to use the standard framework of stochastic thermodynamics SekimotoBook ; Jarzynski11 ; Seifert12 , as we discuss below.
Figure 2: (Color online) The plots of the average velocity $\langle V\rangle$ [see Eq. (24)] and the average entropy production rate $\langle\dot{\sigma}\rangle$ [see Eq. (25)] of the two-sphere microswimmer when the potential is $U(x)=Kx^{2}/2$ and the friction coefficients are $\gamma_{1}(x)=\gamma_{2}(x)=\gamma[1+\delta\tanh(x/w)]$ [see Eq. (22)]. The dimensionless energy parameter is chosen here as $\varepsilon=Kw^{2}/(k_{\mathrm{B}}T_{1})=1$. (a) The plot of the dimensionless average velocity $\langle\hat{V}\rangle=\langle V\rangle\gamma/\sqrt{Kk_{\mathrm{B}}T_{1}}$ as a function of the dimensionless feedback strength parameter $\delta$ for different values of the dimensionless temperature ratio $\theta=T_{2}/T_{1}$. (b) The plot of the dimensionless average entropy production rate $\langle\hat{\dot{\sigma}}\rangle=\langle\dot{\sigma}\rangle\gamma T_{1}/(Kw)^{2}$ as a function of $\delta$ for different values of $\theta$. In both (a) and (b), the filled circles correspond to the result of the numerical simulation.
Figure 3: (Color online) The color plot of the efficiency $\eta$ [see Eq. (26)] of the two-sphere microswimmer as a function of the dimensionless feedback strength parameter $\delta$ and the dimensionless temperature ratio $\theta=T_{2}/T_{1}$. The efficiency $\eta$ vanishes when $\delta=0$.
@@ V Entropy Production Rate
In this section, we calculate the average entropy production rate of the elastic microswimmer in the steady state, which is given by
@@
where the stochastic heat flow $\dot{Q}_{i}$ is the heat gained by the $i$-th sphere per unit time. According to the stochastic energetics developed by Sekimoto SekimotoBook , it is given by
@@
where $\circ$ denotes the Stratonovich product, and the small increment $dx_{i}$ can be obtained from the Langevin equations in Eqs. (5) and (6). Using the Wong-Zakai theorem to deal with the Stratonovich product and taking the statistical average SekimotoBook , we obtain after some calculation
@@
It is straightforward to verify that $\langle\dot{Q}_{i}\rangle=0$ when $\theta=1$ by using the equilibrium distribution function in Eq. (13). When $\theta\neq 1$, we now consider the total heat flow defined by $\langle\dot{Q}\rangle=\langle\dot{Q}_{1}\rangle+\langle\dot{Q}_{2}\rangle$. Using Eqs. (8) and (9), one can generally show that $\langle\dot{Q}\rangle$ vanishes in the steady state because
@@
In the second equality, we have used Eq. (11), which stems from the steady-state condition. Notice that the relation $\langle\dot{Q}\rangle=0$ generally holds even if the two friction coefficients $\gamma_{1}(x)$ and $\gamma_{2}(x)$ are different. Since we have $\langle\dot{Q}_{2}\rangle=-\langle\dot{Q}_{1}\rangle$ from Eq. (20), the average entropy production rate in Eq. (17) can be rewritten as